# Compositional Inversion for Stable Diffusion Models

### 14 Dec 2023 v2

@@ Introduction
Despite the presence of promising outcomes, the composition of inverted concepts with other concepts proves to be challenging. As shown in Fig. LABEL:fig:super_imgs, the results maintain fidelity to the user samples for cat*, but the concept butterflies is absent. This occurs because the method primarily emphasizes the reconstruction loss while disregarding the compositional aspect of the target concept in relation to others. Similar findings are reported in (Tewel et al. 2023) which suggests the dominance of the inverted concepts in the generation process encroaches upon the spotlight of other concepts. However, this is simply attributed to an over-fitting problem, with underlying rationale remaining unexplored.
@@
The larger entropy consequently causes the dominance of the inverted concept over others in the diffusion module. The diffusion module utilizes Transformer blocks to transfer text semantics into visual content, where embeddings are employed to construct the K, and the Q is typically initialized with random noise. Therefore, the presence of a concept heavily relies on the cross-attention of its embedding to the random noise. As the larger entropy of the inverted concept implies a broader span of dimensions to store semantics, it may have a higher probability of obtaining greater attention similarity compared to other concepts. In Fig. 2, we present statistics of the development of the attention similarity between K and Q over iterations. It is evident that the divergence between the inverted concept and others is much more pronounced than that between a pretrained concept and others. The cross-attention mechanism iteratively integrates the attention map of the inverted concept into other concepts, ultimately resulting in the absence of other concepts or their replacement with the inverted concept (e.g., Fig. 2).
@@
Text-to-image synthesis has earned significant attention for its potential applications in content creation, virtual reality, and computer graphics. The objective is to bridge the semantic gap and enable machines to understand and generate images that align with text prompts. For several years, generative adversarial networks (GANs) (Goodfellow et al. 2014; Karras, Laine, and Aila 2019) have been the dominant approach (Zhu et al. 2019; Tao et al. 2022). With recent improvements in DDPM (Ho, Jain, and Abbeel 2020) and DDIM (Song, Meng, and Ermon 2020), text-conditioned diffusion models have made remarkable progress. Building upon the concept of latent images, the Latent Diffusion Model (LDM) (Rombach et al. 2022) was introduced and further extended to Stable Diffusion (Rombach et al. 2022), which is regarded as one of the most promising models for text-to-image synthesis. Another notable framework, Imagen (Saharia et al. 2022), developed by Google, takes a different approach by diffusing pixels directly using a pyramid structure, without relying on latent images. DALL•E2 (Ramesh et al. 2022), developed by OpenAI, uses a prior network that takes text embedding as input to produce an image embedding as the input of the diffusion model. While other alternative models exist (Ramesh et al. 2021; Gu et al. 2022; Meng et al. 2021), we omit further details due to space limitations. Among these models, Stable Diffusion is widely favored by researchers due to its open-source nature.
@@
Current methods in the field of compositionality primarily focus on combining inverted concepts with each other rather than with a broader range of pretrained concepts. This approach, known as multi-concept composition, is related to but distinct from the scope of this paper. Existing methods include Custom Diffusion, SVDiff, Cones, and Perfusion. Custom Diffusion archives this by merging the outputs of multiple models that have been fine-tuned to invert various CoI. It can be considered as a model-level composition approach. SVDiff manually combines objects selected from different CoI concepts as training images, enabling the model to learn to compose them. Cones evaluates the neurons’ contributions to the fidelity of inverted concepts and deactivates those with minor contributions during composition. Perfusion fuses the V components of inverted concepts to balance their contribution to generation. These methods all rely on training/fine-tuning, which requires effort and expertise to gather the regularization images. In contrast, the compositional inversion proposed in this paper is a post-training approach that can be applied to any trained or fine-tuned models and thus is compatible to all the aforementioned methods. Furthermore, the proposed method can be employed for the composition of both pretrained and inverted concepts, making this paper a study on compositionality in a more general sense.
@@
In terms of imposing spatial constraints, there is another category of methods specifically designed for controlling the contours, shapes, or layouts of objects. ControlNet (Zhang, Rao, and Agrawala 2023) trains a new branch that incorporates spatial constraints as input and injects them into each layer of the diffusion module for customized synthesis. Prompt-to-prompt (Hertz et al. 2022) enables object-specific editing by replacing the attention map in the cross-attention module. GLIGEN (Li et al. 2023) designs a gated self-attention layer to incorporate spatial conditions, such as bounding boxes. Layout-control (Chen, Laina, and Vedaldi 2023) employs a training-free approach that ensures higher activation values of the attention maps within the bounding box regions. ReCo (Yang et al. 2023) achieves layout control by encoding regional tokens as part of the text prompt. The spatial inversion module in our proposed method draws inspiration from these methods in terms of controlling the layout. However, these methods are not developed for inversion purpose but rather assume the constrains are known as a prior, while our focus is on automatically discovering the underlying spatial distribution without relying on any user specifications.
@@
where the weight ${\alpha_{t}}$ controls the variance of the Gaussian noises $\mathbf{\epsilon}_{t}$. With the reparameterization trick (Ho, Jain, and Abbeel 2020), Eq. (1) can be written as a close form
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon}_{0},$ (2)
where $\bar{\alpha}_{t}=\Pi_{s=1}^{t}\alpha_{t}$, which makes the forward process deterministic on the initial inputs $\mathbf{x}_{0}$ and a fixed noise $\mathbf{\epsilon}_{0}$. It also defines a Gaussian distribution $q(\mathbf{x}_{t}|\mathbf{x}_{0})$ that we can use to sample latent representations for $\mathbf{x}_{t}$ in the generation.
@@
The learning of the model can be done based on the loss of the estimation $\hat{\mathbf{x}}_{0}$ from ${\mathbf{x}}_{0}$. Modern generative models like Stable Diffusion (Rombach et al. 2022) modify the reverse process to implement guided generations by imposing conditions (e.g., text prompts, depth maps). More specifically, text embedding $\mathbf{e}$ will be fused with $\mathbf{x}_{t}$ to generate a conditioned image using Eq. (3) as $\hat{\mathbf{x}_{0}}=f(\mathbf{x}_{t}\circ\mathbf{e};\mathbf{\theta})$ where $\circ$ is a reserved fusion operator which is implemented differently in various models. The loss is then written
@@
As the $q(\mathbf{x}_{t}|\mathbf{x}_{0})$ in the forward process is a Gaussian distribution, a random vector $\hat{\mathbf{\epsilon}}$ sampled with it can be used as the latent representation for $\mathbf{x}_{t}$. The model can then generate novel images by using these random latent vectors in place of $\mathbf{x}_{t}$ as $\hat{\mathbf{x}}=f(\hat{\mathbf{\epsilon}}\circ\mathbf{e};\mathbf{\theta})$. We can see the generated content is indeed controlled by the only input $\mathbf{e}$. In other words, we can generate desired content as long as we know its text embeddings. This is easy for pretrained concepts (e.g., cat) because the learners have seen enough samples during the training, but hard for specific concepts (e.g., my own cat).
To address this issue, Textual Inversion is a method to backtrack the text embeddings of specific concepts (Gal et al. 2022). It feeds a few samples of the target concept (e.g., 3–5 images of the user’s cat) and a pseudo-concept embedding ($*\mathbf{e}_{\textit{cat*}}$) into the models for inference till the gradients can be calculated and back-propagated to update the pseudo-embedding. It is formulated as

### 11 Jan 2024 v3

@@ Introduction
Figure 1: Image synthesis using traditional inversion methods and the proposed compositional inversion: concepts of ${butter}{flies}$, $street$, and $spaceship$ are absent when composed with concepts inverted with traditional methods.
@@
Despite the presence of promising outcomes, the composition of inverted concepts with other concepts proves to be challenging. As shown in Fig. 1, the results maintain fidelity to the user samples for cat*, but the concept butterflies is absent. This occurs because the method primarily emphasizes the reconstruction loss while disregarding the compositional aspect of the target concept in relation to others. Similar findings are reported in (Tewel et al. 2023) which suggests the dominance of the inverted concepts in the generation process encroaches upon the spotlight of other concepts. However, this is simply attributed to an over-fitting problem, with underlying rationale remaining unexplored.
@@
The larger entropy consequently causes the dominance of the inverted concept over others in the diffusion module. The diffusion module utilizes Transformer blocks to transfer text semantics into visual content, where embeddings are employed to construct the K, and the Q is typically initialized with random noise. Therefore, the presence of a concept heavily relies on the cross-attention of its embedding to the random noise. As the larger entropy of the inverted concept implies a broader span of dimensions to store semantics, it may have a higher probability of obtaining greater attention similarity compared to other concepts. In Fig. 3, we present statistics of the development of the attention similarity between K and Q over iterations. The divergence between the inverted concept and others is much more pronounced than that between a pretrained concept and others. The cross-attention mechanism iteratively integrates the attention map of the inverted concept into other concepts, ultimately resulting in the absence of other concepts or their replacement with the inverted concept (e.g., Fig. 3).
@@
Text-to-image synthesis has earned significant attention for its potential applications in content creation, virtual reality, and computer graphics. The objective is to bridge the semantic gap and enable machines to understand and generate images that align with text prompts. For several years, generative adversarial networks (GANs) (Goodfellow et al. 2014; Karras, Laine, and Aila 2019) have been the dominant approach (Zhu et al. 2019; Tao et al. 2022). With recent improvements in DDPM (Ho, Jain, and Abbeel 2020) and DDIM (Song, Meng, and Ermon 2020), text-conditioned diffusion models have made remarkable progress. Building upon the latent images, the Latent Diffusion Model (LDM) (Rombach et al. 2022) was introduced and further extended to Stable Diffusion (Rombach et al. 2022), which is regarded as one of the most promising models for text-to-image synthesis. Another notable framework, Imagen (Saharia et al. 2022) takes a different approach by diffusing pixels directly using a pyramid structure, without relying on latent images. DALL•E2 (Ramesh et al. 2022) uses a prior network that takes text embedding as input to produce an image embedding as the input of the diffusion model.
@@
Current methods in the field of compositionality primarily focus on combining inverted concepts with each other rather than with a broader range of pretrained concepts. This approach, known as multi-concept composition, is related to but distinct from the scope of this paper. Existing methods include Custom Diffusion, SVDiff, Cones, and Perfusion. Custom Diffusion archives this by merging the outputs of multiple models that have been fine-tuned to invert various CoI. It can be considered as a model-level composition approach. SVDiff manually combines objects selected from different CoI concepts as training images, enabling the model to learn to compose them. Cones evaluates the neurons’ contributions to the fidelity of inverted concepts and deactivates those with minor contributions during composition. Perfusion fuses the V components of inverted concepts to balance their contribution to generation. These methods all rely on training/fine-tuning, which requires effort and expertise to gather the regularization images. In contrast, the compositional inversion proposed in this paper is a post-training approach that can be applied to any trained or fine-tuned models and thus is compatible to all the aforementioned methods. Furthermore, the proposed method can be employed for the composition of both pretrained and inverted concepts, making this paper a study of compositionality in a more general sense.
@@
In terms of imposing spatial constraints, there is another category of methods specifically designed for controlling the contours, shapes, or layouts of objects. ControlNet (Zhang, Rao, and Agrawala 2023) trains a new branch that incorporates spatial constraints as input and injects them into each layer of the diffusion module for customized synthesis. Prompt-to-prompt (Hertz et al. 2022) enables object-specific editing by replacing the attention map in the cross-attention module. GLIGEN (Li et al. 2023) designs a gated self-attention layer to incorporate spatial conditions, such as bounding boxes. Layout-control (Chen, Laina, and Vedaldi 2023) employs a training-free approach that ensures higher activation values of the attention maps within the bounding box regions. ReCo (Yang et al. 2023) achieves layout control by encoding regional tokens as part of the text prompt. The spatial inversion module in our proposed method draws inspiration from these methods in terms of controlling the layout. However, these methods are not developed for inversion purpose but rather assume the constrains as a prior, while our focus is on automatically discovering the underlying spatial distribution without any user specifications.
@@
where ${\alpha_{t}}$ controls the variance of the Gaussian noises $\mathbf{\epsilon}_{t}$. It also defines a Gaussian distribution $q(\mathbf{x}_{t}|\mathbf{x}_{0})$ that we can use to sample latent representations for $\mathbf{x}_{t}$ in the generation.
@@
The learning of the model can be done based on the loss of the estimation $\hat{\mathbf{x}}_{0}$ from ${\mathbf{x}}_{0}$. To implement text-to-image synthesis, text embedding $\mathbf{e}$ will also be fused with $\mathbf{x}_{t}$ to generate a conditioned image using Eq. (2) as $\hat{\mathbf{x}_{0}}=f(\mathbf{x}_{t}\circ\mathbf{e};\mathbf{\theta})$ where $\circ$ is a reserved fusion operator which is implemented differently in various models. The loss is then written
@@
As the generated content is indeed controlled by the only input $\mathbf{e}$, we can generate desired content as long as we know its text embeddings. This is easy for pretrained concepts (e.g., cat) because the learners have seen enough samples during the training, but hard for specific concepts (e.g., my own cat). To address this issue, Textual Inversion is a method to backtrack the text embeddings of specific concepts (Gal et al. 2022). It feeds a few samples of the target concept (e.g., 3–5 images of the user’s cat) and updates the pseudo-concept embedding ($*\mathbf{e}_{\textit{cat*}}$). It is formulated as